OpenStudy (anonymous):

Which lengths could be the sides of a triangle? A. 17 cm, 5 cm, 13 cm B. 5 cm, 17 cm, 12 cm C. 20 cm, 15 cm, 4 cm D. 15 cm, 20 cm, 3 cm

OpenStudy (anonymous):

a

OpenStudy (callisto):

For a triangle, the sum of length of 2 sides must be greater than the other one That is, a+b >c and b+c>a and a+c>b so you should check with groups of length satisfies the above conditions. and that is the answer

OpenStudy (anonymous):

a- b < c

OpenStudy (anonymous):

a+b > c and a- b < c

OpenStudy (anonymous):

basic inquation triangles

OpenStudy (anonymous):

pytago for triangles 90 ; question for lengths

OpenStudy (callisto):

For (A) the three sides are 17 cm, 5 cm, 13 cm 17 +5 = 22 > 13 5+13 = 18 > 17 13 + 17 = 30 > 5 So the three sides can form a triangle For(b), the three sides are 5 cm, 17 cm, 12 cm 5+17 = 22 >12 17 +12 = 29 >5 5+12 = 17 = 17 So, they cannot form a triangle For (c), the three sides are 20 cm, 15 cm, 4 cm 20 +15 = 35 >4 15 + 4 = 19 <20 So, they cannot form a triangle For (d) , the three sides are 15 cm, 20 cm, 3 cm 15 +20 = 35 >3 20 +3 = 23 >15 15 +3 = 18 < 20 So, they cannot form a triangle I hope you can understand how to get the answer...